# Reviewing Scientific Papers for Critical Problems With Reasoning LLMs: Baseline Approaches and Automatic Evaluation

### v2

@@ Abstract
Recent advancements in large language models have sparked interest in utilizing them to aid the peer review process of scientific publication amid the peer review crisis. However, having AI models generate full reviews in the same way as human reviewers risks exacerbating the irresponsible use of LLM-generated reviews. As an alternative, we propose adopting LLMs as manuscript quality checkers. We introduce several baseline approaches and an extendable automatic evaluation framework using top reasoning LLMs as judges to tackle the difficulty of recruiting domain experts for manual evaluation. Utilizing papers withdrawn from arXiv, we validated our proposed methods with several leading reasoning LLMs from multiple vendors and assessed their performance and API costs for identifying critical errors and unsoundness problems in scientific papers. o3 exhibited the best problem identification performance among all models at a modest cost. This paper provides insights into document-based scientific understanding/reasoning and lays a foundation for future applications. Our dataset, code, and model outputs are publicly available.
Reviewing Scientific Papers for Critical Problems With Reasoning LLMs: Baseline Approaches and Automatic Evaluation
Tianmai M. Zhang  Neil F. Abernethy University of Washington {tianmai, neila}@uw.edu
@@ 1 Introduction
Recent advancements in the domain intelligence of large language models (LLMs) have initiated interest in utilizing them to aid the peer review process of scientific publication, especially in consideration of the peer review crisis due to the skyrocketing number of paper submissions in recent years (Kim et al., 2025). Researchers have reported receiving reviews that have likely been written with LLMs, and concern is growing over whether and how AI can be responsibly applied to aid peer review (Naddaf, 2025; Kwon, 2025). The extensive knowledge and high efficiency of LLMs seem promising for streamlining laborious peer review, but irresponsible uses of LLM-generated reviews could significantly undermine trust in the long-established peer review process and publisher credibility.
Several studies have formally explored and evaluated the quality of LLM-generated reviews. An early study by Liang et al. (2024) reported substantial overlap between GPT-generated reviews and human reviews, and participant-reported helpfulness of GPT-generated feedback. Later studies (Du et al., 2024; Zhou et al., 2024; Shin et al., 2025) revealed defects in LLM-generated reviews, such as superficial comments and a lack of criticism or novelty assessment. Other studies (D’Arcy et al., 2024; Gao et al., 2024; Taechoyotin et al., 2024; Tan et al., 2024; Tyser et al., 2024; Yu et al., 2024; Zhu et al., 2025) developed technical methods to improve LLM review generation. However, all of these studies focused on the scenario where LLMs generate full reviews in the same way as human reviewers, which risks exacerbating the irresponsible use of LLM-generated reviews. The most common evaluation method in these studies was comparing LLM-generated reviews with human reviews, either manually or computationally.
@@ 2.1 Dataset
Since retraction comments were typically short and did not necessarily contain clear mentions of errors (e.g., "This paper has been withdrawn due to some mistakes"), we further filtered the dataset with the help of LLMs. Specifically, de-identified retraction comments were first provided to Gemini 2.5 Flash (preview-04-17) to determine whether each retraction comment clearly specified the error. This step resulted in a subset of 2,190 cases. Our manual review further excluded cases that (1) were incorrectly identified during LLM screening; (2) belonged to different versions of the same paper; (3) were not in English; (4) suspiciously used exactly the same retraction reason ("The author has withdrawn this paper due to a crucial sign error in equation 1") that seems to have been a template provided by arXiv in early years; (5) contained problems that were unlikely to be detectable from the manuscript alone (e.g., an error in a core reference, a bug in code, unreliable raw data, problems revealed by new observations, etc.). We also corrected mistakenly redacted theorem names in the retraction comments. The final dataset, named WithdrarXiv-Check, contains 1,225 cases in total.
@@ 2.4 Experiment Setup
Table 2: Numbers of identified problems, performance metrics, token usage, and estimated costs of LLM checkers.
3 Results
Table 2 provides the main results assessing LLM quality checkers. Besides numbers of identified problems and performance metrics, we recorded token usage to inform future work that seeks to apply LLMs at a larger scale. Average costs of reviewing a paper under each pipeline-LLM combination were estimated based on the standard API pricing [^6] On June 10, OpenAI dropped the price of o3 API by 80%. of the LLM vendors in June 2025.
@@ 3 Results
Additional evaluation results by individual judges are shown in Table 3 in the Appendix. Compared to performance scores determined by a single judge, final scores reasonably dropped due to the difficulty of receiving affirmative votes from both judges, which demonstrates some resistance of our multi-judge approach to potential false positives in automatic evaluation. The Gemini 2.5 Pro judge is consistently more lenient than o3, sometimes making hallucinatory assumptions about the relationship between a problem submission and the gold error description. For example, when comparing an LLM checker’s report of a missing factor in an unnumbered equation after a specific sentence and a retraction comment regarding an error in equation 13, the Gemini 2.5 Pro judge said it is "extremely likely" that the erroneous equation is equation 13. In our tests with different judge instructions, Gemini 2.5 Pro was also far less responsive than o3 to important instructions such as "Default your answer to ’No’ and only give ’Yes’ if you are certain". The phrase "exactly the same problem" in the prompt for judges to determine hits is particularly necessary for Gemini 2.5 Pro, without which it would give an even higher proportion of affirmative votes, although the trend in hit rate results remained the same. These observations reveal preference differences between models from different vendors and thus highlight the necessity of adopting multiple judges or model settings instead of a single judge while using the LLM-as-a-judge approach.
@@ 4 Discussion
In this work, we only assessed the performance of the most simplistic approach—providing an LLM with a paper and tasking it with finding problems. More complex approaches are likely to demonstrate further gains in LLM performance. For example, one can further customize the prompt (e.g., by considering the scientific field a paper belongs to), design a propose-then-verify workflow, expand task input (e.g., by including supplementary materials and references of a paper, or additional domain knowledge through retrieval from a knowledge base) and tools (e.g., code execution), fine-tune models for specific reviewing tasks, or adopt multi-agent collaboration.
@@ 5 Limitation
As a proof-of-concept study, this work has several limitations. First, only closed-source reasoning LLMs were evaluated. Future work may consider comparing different PDF preprocessing pipelines and open-source LLMs. Second, our evaluation metrics may suffer inaccuracies due to (1) ambiguities in paper authors’ retraction comments (e.g., "withdrawn due to a crucial error in Lemma 2", without specification of the error, (2) automatic evaluation utilizing LLMs without human involvement, and (3) our settings that each LLM checker was tested only once per paper and each LLM judge graded each submission only once. Nevertheless, a higher score under parallel evaluation still reasonably indicates better performance of an LLM checker. We expect gradual improvements in the reliability of the LLMs-as-judges approach as leading models continue to evolve. Third, the impacts of experiment parameters or alternate prompts were not formally investigated or optimized. In addition, our results based on a dataset rich in math and physics papers published in the past may not generalize well to papers in other scientific domains or future papers. Last, since we were unable to recruit domain experts for manual annotation, it remains possible that some cases in the dataset contain problems that could not be detected based on the manuscript alone, or that the authors gave inaccurate reasons for retraction, which may pose an upper limit for model performance. If conditions permit, future work should also consider involving domain experts in calibrating the LLM judges to resolve potential biases caused by circular evaluation of LLMs using LLMs.
@@
  • They treated all LLM-identified problems that did not directly match the gold error description as wrong answers, whereas we allowed some flexibility through LLM reasoning and attempted to assess the precision of these additional answers.
@@
Our WithdrarXiv-Check dataset, experiment code, and model outputs (including thinking outputs if available) are available on Github [^10] https://github.com/tianmai-zhang/WithdrarXiv-Check. Readers may use the dataset for further experiments, reproduce the numbers in our tables, or perform a closer inspection of the model outputs to obtain further insights into LLM behaviors. We welcome discussions, collaborations, and funding for future work.
Acknowledgments
We thank community members of the Black Spatula Project for their inspiring discussions, especially Ugo Dos Reis for highlighting the WithdrarXiv dataset in December 2024, Delip Rao for instructions on how to use this dataset, and Casey Wimsatt and David Reinstein for constructive feedback on early versions of this paper. The authors also thank Yuxin Xu for her help with the API.

### v3

# AI for Science

@@ Abstract
Recent advancements in large language models have sparked interest in utilizing them to aid the peer review process of scientific publication amid the peer review crisis. However, having AI models generate full reviews in the same way as human reviewers risks exacerbating the irresponsible use of LLM-generated reviews and instigating intentional manipulation. As an alternative, we propose adopting LLMs as manuscript quality checkers. We introduce several baseline approaches and an extendable automatic evaluation framework using top reasoning LLMs as judges to tackle the difficulty of recruiting domain experts for manual evaluation. Utilizing papers withdrawn from arXiv, we validated our proposed methods with several leading reasoning LLMs available in May-June 2025 and assessed their performance and API costs for identifying critical errors and unsoundness problems in scientific papers. o3 exhibited the best problem identification performance among all models at a modest cost. This paper provides insights into document-based scientific understanding/reasoning and lays a foundation for future applications. Our dataset, code, and model outputs are publicly available.
@@ 1 Introduction
Recent advancements in the domain intelligence of large language models (LLMs) have fostered interest in utilizing them to aid the peer review process of scientific publication, especially in consideration of the peer review crisis due to the skyrocketing number of paper submissions in recent years (Kim et al., 2025). Researchers have reported receiving reviews that have likely been written with LLMs, and concern is growing over whether and how AI can be responsibly applied to aid peer review (Naddaf, 2025; Kwon, 2025). The extensive knowledge and high efficiency of LLMs seem promising for streamlining laborious peer review, but irresponsible uses of LLM-generated reviews could significantly undermine trust in the long-established peer review process and publisher credibility.
Several studies have formally explored and evaluated the quality of LLM-generated reviews. An early study by Liang et al. (2024) reported substantial overlap between GPT-generated reviews and human reviews, as well as participant-reported helpfulness of GPT-generated feedback. Later studies (Du et al., 2024; Zhou et al., 2024; Shin et al., 2025) revealed defects in LLM-generated reviews, such as superficial comments and the lack of criticism or novelty assessment. Other studies (D’Arcy et al., 2024; Gao et al., 2024; Taechoyotin et al., 2024; Tan et al., 2024; Tyser et al., 2024; Yu et al., 2024; Zhu et al., 2025; Zeng et al., 2025) developed technical methods to improve LLM review generation. However, all of these studies focused on the scenario where LLMs generate full reviews in the same way as human reviewers, which risks exacerbating the irresponsible use of LLM-generated reviews and instigating intentional manipulation (Ye et al., 2024; Lin, 2025). The most common evaluation method in these studies was comparing LLM-generated reviews with human reviews, either manually or computationally.
@@ 2.1 Dataset
Since retraction comments were typically short and did not necessarily contain clear mentions of errors (e.g., "This paper has been withdrawn due to some mistakes"), we further filtered the dataset with the help of LLMs. Specifically, de-identified retraction comments were first provided to Gemini 2.5 Flash (preview-04-17) to determine whether each retraction comment clearly specified the error. This step resulted in a subset of 2,190 cases. Our manual review further excluded cases that (1) were incorrectly identified during LLM screening; (2) belonged to different versions of the same paper; (3) were not in English; (4) suspiciously used exactly the same retraction reason (e.g., "The author has withdrawn this paper due to a crucial sign error in equation 1") that seems to have been a template provided by arXiv in early years; (5) contained problems that were unlikely to be detectable from the manuscript alone (e.g., an error in a core reference, a bug in code, unreliable raw data, problems revealed by new observations, etc.). We also corrected mistakenly redacted theorem names in the retraction comments. The final dataset, named WithdrarXiv-Check, contains 1,225 cases in total.
@@ 2.4 Experimental Setup
3 Results
Table 2 provides the main results assessing LLM quality checkers. Besides numbers of identified problems and performance metrics, we recorded token usage to inform future work that seeks to apply LLMs at a larger scale. Average costs of reviewing a paper under each pipeline-LLM combination were estimated based on the standard API pricing of the LLM vendors in June 2025.
Table 2: Numbers of reported problems, performance, token usage, and costs of LLM checkers
@@ 3 Results
Additional evaluation results by individual judges are shown in Table 3. Compared to performance scores determined by a single judge, final scores reasonably dropped due to the difficulty of receiving affirmative votes from both judges, which demonstrates some resistance of our multi-judge approach to potential false positives in automatic evaluation. The Gemini 2.5 Pro judge is consistently more lenient than o3, sometimes making hallucinatory assumptions about the relationship between a problem submission and the gold error description. For example, when comparing an LLM checker’s report of a missing factor in an unnumbered equation after a specific sentence and a retraction comment regarding an error in equation 13, the Gemini 2.5 Pro judge said it is "extremely likely" that the erroneous equation is equation 13. In our tests with different judge instructions, Gemini 2.5 Pro was also far less responsive than o3 to important instructions such as "Default your answer to ‘No’ and only give ‘Yes’ if you are certain". The phrase "exactly the same problem" in the prompt for judges to determine hits is particularly necessary for Gemini 2.5 Pro, without which it would give an even higher proportion of affirmative votes, although the trend in hit rate results remained the same. These observations reveal preference differences between models from different vendors and thus highlight the necessity of adopting multiple judges or model settings instead of a single judge while using the LLM-as-a-judge approach.
Table 3: Evaluation results by individual judges
Hit Rate Evaluation
Checker Model Judge Model
Gemini 2.5 Pro o3 (medium)
PDF as an attachment
Gemini 2.5 Pro 68.2 41.2
Gemini 2.5 Flash 66.5 40.4
o3 (medium) 80.4 50.2
o4-mini (medium) 72.7 40.4
Claude 3.7 Sonnet 24.1 12.2
LaTeX script in prompt
Gemini 2.5 Pro 64.5 38.0
Gemini 2.5 Flash 64.1 35.9
o3 (medium) 79.2 51.4
o4-mini (medium) 71.8 40.0
Claude 3.7 Sonnet 48.6 26.5
Average Precision Evaluation
Checker Model Judge Model
Gemini 2.5 Pro o3 (medium)
PDF as an attachment
Gemini 2.5 Pro 70.7 39.7
Gemini 2.5 Flash 46.4 29.8
o3 (medium) 55.0 36.9
o4-mini (medium) 48.3 32.7
Claude 3.7 Sonnet 51.1 47.5
@@ 4 Discussion
In this work, we only assessed the performance of the most simplistic approach—providing an LLM with a paper and tasking it with finding problems. More complex approaches or newer generations of models are likely to demonstrate further gains in LLM performance. Examples of these approaches include further customizing the prompt (e.g., by considering the scientific field a paper belongs to), designing a propose-then-verify workflow, expanding task input (e.g., including supplementary materials, paper references, or additional domain knowledge) and tools (e.g., code execution), fine-tuning models for specific reviewing tasks, or adopting multi-agent collaboration.
@@ 5 Limitation
As a proof-of-concept study, this work has several limitations. First, only closed-source reasoning LLMs were evaluated. Future work may consider comparing different PDF preprocessing pipelines and open-source LLMs. Second, our evaluation metrics may suffer inaccuracies due to (1) ambiguities in paper authors’ retraction comments (e.g., "withdrawn due to a crucial error in Lemma 2", without further specification of the error, (2) automatic evaluation utilizing LLMs without human involvement, and (3) our settings that each LLM checker was tested only once per paper and each LLM judge graded each submission only once. Nevertheless, a higher score under parallel evaluation still reasonably indicates better performance of an LLM checker. We expect gradual improvements in the reliability of the LLMs-as-judges approach as leading models continue to evolve. Third, the impacts of experiment parameters or alternate prompts were not formally investigated or optimized. In addition, our results based on a dataset rich in math and physics papers published in the past may not generalize well to papers in other scientific domains or future papers. Last, since we were unable to recruit domain experts for manual annotation, it remains possible that some cases in the dataset contain problems that could not be detected based on the manuscript alone, or that the authors gave inaccurate reasons for retraction, which may pose an upper limit for measurement of model performance using our dataset. If conditions permit, future work should also consider involving domain experts in calibration of LLM judges to resolve potential biases caused by circular evaluation of LLMs using LLMs.
@@
  • They treated all LLM-identified problems that did not strictly match the gold error description as wrong answers, whereas we allowed some flexibility through LLM reasoning and attempted to assess the precision of these additional answers.
@@
Our WithdrarXiv-Check dataset, experiment code, and model outputs (including thinking outputs if available) are available on Github [^9] https://github.com/tianmai-zhang/WithdrarXiv-Check. Readers may use the dataset for further experiments, reproduce the numbers in our tables, or perform a closer inspection of the model outputs to obtain further insights into LLM behaviors.
Acknowledgments and Disclosure of Funding
This work is supported by the Ira Kalet and Fred Wolf Endowment Fund from the Department of Biomedical Informatics and Medical Education, University of Washington. The authors declare no competing interest.
The authors thank community members of the Black Spatula Project for their inspiring discussions, especially Ugo Dos Reis for highlighting the WithdrarXiv dataset in December 2024, Delip Rao for instructions on how to use this dataset, and Casey Wimsatt and David Reinstein for constructive feedback on early versions of this paper. The authors also thank Yuxin Xu for her help with the API.